In terms of pie, how much further does a 70cm diameter bicycle tire travel in one revolution than a 50cm bicycle tire?

Answer by longjonsilver(2297)

circumference, C, of the wheel is the distance travelled in 1 revolution, and C=pi*d (or C=2pi*r)
70cm diameter wheel, C=70*pi
50cm diameter wheel, C=50*pi
So, the larger wheel will have travelled 70*pi - 50*pi = (20*pi) cm

jon